OpenStudy (anonymous):

write an equation for the line tangent to f(x) = xe^-x + 1 at c = 0

OpenStudy (anonymous):

i know that the derivative is -xe^-x + e^-x and that f '(0) = 1. but now what?

OpenStudy (anonymous):

calculate f(0) it'll be 1 also then write de tangecial equation like : y=f(c)+f'(c)(x-c) were c=0 it should look like this : y=1+1(x-0) (=) y=1+x hope this helps

OpenStudy (anonymous):

thanks!!